19a-3(n+3)=9 whats n??

Answer by ankor@dixie-net.com(12692)

19a - 3(n+3) = 9; what's n??
:
Mult what's inside the brackets:
19a - 3n - 3 = 9
19a - 3n = 9 + 3
19a - 3n = 12
-3n = 12 - 19a
:
Divide both sides by -3
n = (12/-3) - (19a/-3)
n = -4 - (-19a/3)
n = -4 + (19a/3)
or
n = (19a/3) - 4